Find the length of an arc in a circle of radius 10 centimeters subtended by the central angle of 50°. Show your work.


2. Graph on [-4π, 4π] and verbalize how the graph varies from the graphs of . 

Graph  on the window [−5π, 5π] and describe freely what the graph shows. You can use www.desmos.com/calculator to obtain the graphs.


3. A 23-ft ladder leans against a building so that the angle between the ground and the ladder is 80°. How high does the ladder reach up the side of the building? Show the steps of your reasoning.  


1
Expert's answer
2021-12-30T01:19:39-0500

1. The circumference of the circle of radius "r" is "L=2\pi r."

An arc is a portion of the outline of a circle.

The length of the arc of a circle of radius "r" that subtends an angle of "\u03b8"

 in degrees at the center is


"Arc\ length=2\pi r(\dfrac{\theta}{360\degree})"

Given "r=10\ cm, \theta=50\degree."


"Arc\ length=2\pi (10)(\dfrac{50\degree}{360\degree})=\dfrac{25\pi}{9}(cm)"

"=\dfrac{25\pi}{9}(cm)\approx8.7266(cm)"

2.

(i)

"f(x)=x\sin x"

First draw the graph of "f(x)=\u00b1x."

The function "f(x)=x\sin x" is even. The graph is symmetric with respect to the "y" -axis.

Points "(-\dfrac{\pi}{2}+2\pi n,\dfrac{\pi}{2}+2\pi n),n\in \Z" lie on the graph of "f(x)=-x"


Points "(\dfrac{\pi}{2}+2\pi n,\dfrac{\pi}{2}+2\pi n),n\in \Z" lie on the graph of "f(x)=x"


Points "(-\dfrac{3\pi}{2}+2\pi n,-\dfrac{3\pi}{2}+2\pi n),n\in \Z" lie on the graph of "f(x)=x"


Points "(\dfrac{3\pi}{2}+2\pi n,-\dfrac{3\pi}{2}+2\pi n),n\in \Z" lie on the graph of "f(x)=-x"




(ii)

The function "f(x)=\dfrac{\sin x}{x}" is not defined at "x=0."


"\lim\limits_{x\to 0}f(x)=\lim\limits_{x\to 0}\dfrac{\sin x}{x}=1"

The function "f(x)=\dfrac{\sin x}{x}" has a removable discontinuity at "x=0."

The function "f(x)=\dfrac{\sin x}{x}" is even. The graph is symmetric with respect to the "y" -axis.

"y\to0" as "x\to\pm \infin."



The graph of "f(x)=\dfrac{\sin x}{x}" is decaying oscillations (oscillations of continuously decreasing amplitude). The oscillations never stop, but go on decreasing in strength. The amplitude of oscillations at any point "x\not=0" is "1\/x."


3.



From right triangle


"h=L\sin \theta"

"h=23\sin 80\degree \ ft\approx22.65\ ft"